# Unsupervised Text Embedding Space Generation Using Generative Adversarial Networks for Text Synthesis

### v1

@@ 3.3 Generator
Both real and fake seeds ($H_{real}$ and $H_{fake}$) are essential for adversarial training. The real seeds can be obtained from the real text as described in the seed interpretation model, and the fake seeds are generated by the generator. In most text-GAN models reported so far, fake sentences are obtained from a text-based pre-trained autoregressive generator. Thus, data memorization that several synthetic sentences reproduce the training data occurs. To prevent data memorization, our generator does not use a pre-trained autoregressive-based model and does not explicitly refer to the text in the training data during adversarial training. Our generator aims to create suitable fake text embedding spaces (fake seeds, $H_{fake}$) in an unsupervised way (GAN framework) by referencing continuous spaces (real seeds). The embedding space created by the generator has the same dimension of (sequence length * embedding dimension), just as the real seed.
The generator $g_{\phi}(\cdot)$ is composed of two convolutional layers and generates seeds from the uniform distribution noise $X$ with an interval of $[-10,10)$ to make diverse forms of the seeds. Also, using random noise has the advantage of not having to select the first token in the text synthesis process after model learning. The final $H_{fake}$ can be obtained by Equation 6:
$\small H_{fake}=g_{\phi}\big{(}X\sim U(-10,10)\big{)}$ (6)
Moreover, we compare one more additional model, perturbed TESGAN (P-TESGAN). P-TESGAN creates perturbed seeds by adding zero-centered normal distribution noise $z$ to the generator output. P-TESGAN is expected to learn more robustly by perturbating the generator output. Further details of the generator are explained in Appendix A.
@@ 3.4 Objective Functions
Since the generator does not refer to text while adversarial training, its performance is determined by the loss of the discriminator. Thus, we propose four types of loss to update the parameters of the generator and the discriminator.
@@ 3.4.1 Discriminators
Sentence structure is important for constructing a complete sentence. Since the real seeds are made from perfect sentences, they maintain structural representations of sentences. Therefore, the fake seeds should capture the structural features of the real seeds. We use Bidirectional Encoder Representations from Transformer (BERT) [6] $d_{\alpha}(\cdot)$, which called Seed Structure Discriminator (SSD), to capture the structural features of sentences, and the first token is used to predict whether the seed is real or fake.
The order of tokens is also important for constructing sentences. It is possible to predict whether a sentence order representation of a seed is correct because both real and fake seeds have a dimension of (sequence length * embedding dimension). To do this, we use Bidirectional LSTM $d_{\beta}$, which called Seed Order Discriminator (SOD), to consider both forward and backward directions of sentences. Concatenated the first and the last hidden states are used to predict whether the seed is real or fake. Finally, SSD and SOD updates both the discriminator and the generator during adversarial training and their loss function can be defined as the following equation:
$\small\begin{gathered}\mathcal{L}_{D}=-\frac{1}{N}\Sigma_{i=1}^{N}\big{[}y_{i}\log{x_{i}}+(1-y_{i})\log{(1-x_{i})}\big{]}\\ x=predicted,y=target\end{gathered}$ (7)
The detailed specification and figure of the discriminators can be found in Appendix A.
@@ 3.4.2 Generator Helpers
Since capturing the distribution of text embedding space is important, we conduct Seed Distribution Prediction (SDP). However, since the text embedding space cannot be directly used as a probability distribution, the seed interpretation model output is used. Specifically, if a seed passes through the frozen seed interpretation model, the output dimension of (sequence length * vocab size) is obtained by the softmax function which can be used as a probability distribution. As shown in Figure of Appendix A, the loss is calculated by Kullback-Leibler divergence with distributions of the real and the fake seeds. SDP is used only for updating the generator during adversarial training:
$\small\mathcal{L}_{SDP}=\sigma\big{(}f_{\theta}(H_{real})\big{)}\log{\frac{\sigma\big{(}f_{\theta}(H_{real})\big{)}}{\sigma\big{(}f_{\theta}(H_{fake})\big{)}}}$ (8)
where the $\sigma$ and $f_{\theta}(\cdot)$ mean softmax function and the frozen seed interpretation model.
The sentences used as the seeds are composed of tokens explained in Equation 2. Since the structures of seeds are somewhat formalized, we conduct Seed Frame Prediction (SFP). Therefore, we calculate the Mean Absolute Error (MAE) and Mean Squared Error (MSE) to make a form of a fake seed similar to a real one. If we train the fake seeds through the MAE and the MSE, the fake seeds from the generator can be blurred. However, the loss of SFP is quite small compared to that of SSD, SOD, and SDP; therefore, SFP does not harm the generator. SFP is used only for updating the generator during adversarial training:
@@ 4.1 Dataset
Table 1: Statistics of DailyDialog
Data Train Validation Test
Dialogues 11,118 1,000 1,000
Total sentences 87,170 8,069 7,740
Avg. sentences 7.84 8.07 7.74
TESGAN synthesizes text based on the seed obtained from the prior sentence. Therefore, the seed interpretation model should be trained with multi-turn sentence before adversarial training. In this experiment, we use DailyDialog [^1] http://yanran.li/dailydialog, data are submitted in the supplementary. [17] multi-turn conversation data, which is a dataset for training an open-domain dialogue generation model. DailyDialog consists of chit-chat-style multi-turn English conversations, and we select this data for domain-independent text synthesis. Statistics of DailyDialog are shown in Table 1.
@@ 4.2 Training Steps
First, the seed interpretation model must be pre-trained by multi-turn conversation data $D=S_{1},...,S_{N}$ to interpret the seeds. As shown in Table 1, training is conducted with 11k multi-turn dialogue data, and the model achieving the highest BLEU-4 score in the validation set is selected. After pre-training the seed interpretation model, the generator and the discriminator learn through adversarial training. For adversarial training, real and fake seeds are made by the embedding part of the frozen seed interpretation model and the generator, respectively. Since the real seeds can be created from any sentence, all 87k sentences in the training set are used to make the real seeds via Equation 2. We also generate the same number of fake seeds as the real seeds to conduct adversarial training:
@@
where $d_{\alpha,\beta}$ means SSD, SOD respectively. The sentences used as the seeds are composed of tokens explained in Equation 2. Detailed explanation of above pseudocode, optimization methods, and hyperparameters are explained in Appendix B.
@@ 4.3 Evaluation Metric
Target-oriented evaluation metrics, such as BLEU and ROUGE, are not suitable for evaluating synthetic text. This is because each synthesized sentence from random noise has no corresponding target and the generative models aim to synthesize plausible data based on real data distribution without copying the training data. Several methods were suggested that can be used in this case [20, 4].
@@ 4.3.1 Fréchet BERT Distance (FBD)
de Masson d’Autume et al. (2019) [20] proposed Fréchet Embedding Distance (FED) to evaluate the quality of synthetic text inspired by Fréchet Inception Distance (FID) [10]. Alihosseini et al. (2019) [1] proposed FBD, an improved version of FED, to measure the quality and diversity of synthesized text using pre-trained BERT. The features of real and synthesized text obtained by the pre-trained BERT are assumed to have Gaussian distributions and FBD is the distance between them:
@@ 4.3.2 Multi-Sets-Jaccard (MSJ)
Each synthesized sentence has no corresponding target; thus, we select MSJ that calculates the score between real and synthesized text sets. The Jaccard Index determines the similarity of two sets, calculating the ratio of the cardinality of their intersection to that of their union. Inspired by Jaccard Index, MSJ focuses on the similarity of the n-grams frequencies of text in the two sets of $s_{r}$ and $s_{f}$ which are real and synthesized text sets respectively:
$\small MSJ_{n}=\frac{\Sigma_{g\in G_{n}}\min\big{(}C_{n}(g,s_{r}),C_{n}(g,s_{f})\big{)}}{\Sigma_{g\in G_{n}}\max\big{(}C_{n}(g,s_{r}),C_{n}(g,s_{f})\big{)}}$ (12)
where $G_{n}$ and $C_{n}(g,s)$ shows the n-grams in $s_{r}\cup s_{f}$ and the normalized counts of the n-gram in set $s$. Thus MSJ can consider both quality and diversity of synthetic results.
@@ 4.3.3 Language Model score (LM)
de Masson d’Autume et al. (2019) and Caccia et al. (2020) [20, 4] proposed LM which can evaluate the quality of generated samples using a well-trained language model. LM measures generated sample quality which means that scores of the bad samples are poor under a well-trained language model. We select pre-trained GPT-$2$https://huggingface.co/docs/transformers/model_doc/gpt2 as a well-trained language model. LM is calculated as the cross-entropy value of the output and input of GPT-2.
@@ 4.3.4 Data Synthesis Ratio (DSR)
DSR considers not only data memorization ratio between training and synthesized data but also synthetic diversity itself. Short sentences identical to training data, such as "I’m fine", can be synthesized by coincidence. Therefore, sentences longer than two-thirds of the maximum sentence length that perfectly reproduce the training data are considered memorized data. Considering these conditions, we can calculate DSR using the following equation:
@@
where $S_{syn}$ and $S_{train}$ indicate synthesized and training text set respectively. $S_{unq}$ means the set of unique sentences of synthesized text results. If the synthesized sentence in $S_{syn}$ do not reproduce any of the sentences in $S_{train}$, $R_{syn}$ would be 1. Similarly, if the synthesized text in $S_{syn}$ is all unique, $R_{unq}$ will be 1. The final DSR is calculated as the harmonic mean of $R_{syn}$ and $R_{unq}$ ratio.
@@ 4.4 Baselines
Here, we compare our two models with the following approaches: Maximum Likelihood Estimation (MLE), PG-BLEU, SeqGAN, RankGAN, and MaliGAN. MLE is the pre-training result of the generator, which all the other models do before adversarial training. Since MLE is trained without adversarial training, it is shown as a baseline in Figure 2. The pre-trained generators with the lowest loss in the validation set were chosen for each method including MLE. We compare these models with our original TESGAN and P-TESGAN trained by adding zero-centered normal distribution $z$ to the generator output. We also present the results of using Gaussian random noise as input of the seed interpretation model to prove that the results of the TESGAN-based model are due to seeds created by the generator, not the seed interpretation model.
@@ 5 Results
Figure 2: Illustration showing the results of the text-GAN models. In the previous research, adversarial training is conducted after the generator pre-training. MLE is represented as a baseline because it is a supervised pre-trained generator without adversarial training.
Table 2: Performance of the models. P-TESGAN denotes the perturbed TESGAN. Numbers in parentheses indicate the training epoch of the selected model. $\dagger$ is the result of directly entering Gaussian random noise as an input to the seed interpretation model. $\ast$ denotes a metric not considered when selecting the best model. - denotes that the confidence of the result is low because the quality of the synthesized sentence is poor.
Method FBD $\downarrow$ MSJ2 $\uparrow$ MSJ3 $\uparrow$ MSJ4 $\uparrow$ MSJ5 $\uparrow$ DSR ($R_{syn},R_{unq}$) normal-$\uparrow$ LMnormal-$\ast$ $\downarrow$
TESGAN (17) 2.899 0.150 0.079 0.042 0.021 0.967 (1, 0.936) 4.236
P-TESGAN (10) 2.274 0.131 0.066 0.032 0.014 0.841 (0.997, 0.727) 3.642
SeqGAN (1) 6.153 0.185 0.091 0.040 0.015 0.880 (0.883, 0.877) 5.094
RankGAN (1) 6.409 0.189 0.096 0.048 0.023 0.890 (0.895, 0.886) 5.123
MaliGAN (1) 21.436 0.015 0.006 0.003 0 0.030 (1, 0.015) -
PG-BLEU (1) 9.002 0.071 0.036 0.015 0.006 0.569 (0.555, 0.584) 4.584
MLE 6.284 0.193 0.096 0.045 0.021 0.955 (0.925, 0.987) 5.168
Random Noise $\dagger$ 14.142 0.080 0.036 0.016 0.006 0.930 (1, 0.869) 4.562
Table 3: Example of the synthesized sentences. P-TESGAN denotes the perturbed TESGAN.
TESGAN (17-epoch) P-TESGAN (10-epoch) Random Noise

### v4

@@ 3.3 Generator
Figure 3: Illustration of the generator. P-TESGAN makes perturbed seeds by adding zero-centered normal distribution noise $z$ (gray) to the output (blue) from the generator.
Both real and fake seeds ($H_{real}$ and $H_{fake}$) are essential for adversarial training. Real seeds can be obtained from the real text, as described in the seed interpretation model, and fake seeds are generated by the generator. In most text-GAN models reported so far, fake sentences are obtained from a text-based pre-trained autoregressive generator. Consequently, data memorization occurs, where several synthetic sentences reproduce the training data. To prevent data memorization, our generator does not use a pre-trained autoregressive-based model and does not explicitly reference the text in the training data during adversarial training. Our generator aims to create suitable fake text embedding spaces $H_{fake}$ in an unsupervised manner (GAN framework) by referencing real text continuous spaces $H_{real}$.
As shown in Figure 3, the generator $g_{\phi}(\cdot)$ is composed of two convolutional layers and generates seeds from the uniform distribution noise $X$ within an interval of $[-10,10)$ to create diverse forms of the seeds. Additionally, using random noise has the advantage of not having to select the first token in the text synthesis process after model learning. The final $H_{fake}$ can be obtained by Equation 7:
$\small H_{fake}=g_{\phi}\big{(}X\sim U(-10,10)\big{)}\in\mathbb{R}^{L\times d}$ (7)
where $L$ and $d$ represent sequence length and embedding dimensions. As a result, the embedding space created by the generator has the same dimension as the real seed. Moreover, we also compare an additional model, perturbed TESGAN (P-TESGAN). P-TESGAN creates perturbed seeds by adding zero-centered normal distribution noise $z$ to the generator output. P-TESGAN is expected to learn more robustly by perturbating the generator output. Please refer to Appendix A for detailed model information.
@@ 3.4 Objective Functions
Since the generator does not refer to text during adversarial training, its performance is determined by the loss of the discriminator. Thus, we propose four types of loss to update the parameters of the generator and the discriminator.
@@ 3.4.1 Discriminators
(a) Seed Structure Discriminator (SSD)
(b) Seed Order Discriminator (SOD)
Figure 4: Illustrations of the two discriminators. SSD predicts whether the seed is real or fake using the $[CLS]$ special token’s feature. SOD considers both forward and backward contexts of the seed.
Sentence structure is important for constructing a complete sentence. Since the real seeds are made from perfect sentences, they maintain structural representations of sentences. Therefore, the fake seeds should capture the structural features of the real seeds. As shown in Figure 3(a), we use Bidirectional Encoder Representations from Transformer (BERT) Devlin et al. (2019) $d_{\alpha}(\cdot)$ called Seed Structure Discriminator (SSD) to capture the structural features of sentences, and the first hidden state is used to predict whether the seed is real or fake.
The order of tokens is also important for constructing sentences. It is possible to predict whether a sentence’s order representation of a seed is correct because both real and fake seeds have a dimension of (sequence length * embedding dimensions). To do this, as shown in Figure 3(b), we use Bidirectional LSTM $d_{\beta}(\cdot)$ called Seed Order Discriminator (SOD) to consider both forward and backward directions of sentences. The concatenation of the first and the last hidden states is used to predict whether the seed is real or fake.
During adversarial training, both discriminators are trained to predict whether the seeds are real (label 1) or fake (label 0), while the generator is trained to fool the discriminators by predicting fake seeds as 1. The loss function of the discriminators is defined by the following equation, which updates both the discriminators and the generator:
$\small\begin{gathered}\mathcal{L}_{D}=-\frac{1}{N}\sum_{i=1}^{N}\big{[}y_{i}\log{x_{i}}+(1-y_{i})\log{(1-x_{i})}\big{]}\\ x=predicted,y=target\end{gathered}$ (8)
Additional information regarding the size and descriptions of the discriminators can be found in Appendix A.
@@ 3.4.2 Generator Helpers
Figure 5: Illustration of Seed Distribution Prediction (SDP). SDP is used to enhance the fake seeds of the generator during adversarial training by minimizing the distance between real and fake seed distributions.
During adversarial training, it is challenging for the generator to learn solely from the discriminators introduced in Section 3.4.1. Therefore, in this section, two auxiliary tasks are introduced to aid the training of the generator.
Capturing the distribution of the text embedding space is important, and for this purpose, we employ Seed Distribution Prediction (SDP). However, since the text embedding space cannot be directly used as a probability distribution, the output of the seed interpretation model is utilized. Specifically, when a seed passes through the frozen seed interpretation model, the output dimension of (sequence length * vocabulary size) is obtained through the softmax function, which can be used as a probability distribution. The loss is calculated using the Kullback-Leibler divergence between the distributions of the real and the fake seeds. SDP is used solely for updating the generator during adversarial training:
$\small\mathcal{L}_{SDP}=\sigma\big{(}f_{\theta}^{*}(H_{real})\big{)}\log{\frac{\sigma\big{(}f_{\theta}^{*}(H_{real})\big{)}}{\sigma\big{(}f_{\theta}^{*}(H_{fake})\big{)}}}$ (9)
where the $\sigma$ and $f_{\theta}^{*}(\cdot)$ mean softmax function and the frozen seed interpretation model. More detailed figure of SDP is illustrated in Figure 5.
The sentences used as the seeds are composed of tokens explained in Equation 3. Additionally, we apply Seed Frame Prediction (SFP) since the structures of seeds are somewhat formalized. Therefore, we calculate the Mean Absolute Error (MAE) and Mean Squared Error (MSE) to make the form of a fake seed similar to a real one. If we train the fake seeds using MAE and MSE, the fake seeds from the generator can become blurred. However, the loss of SFP is relatively small compared to that of SSD, SOD, and SDP; therefore, SFP does not adversely affect the generator. SFP is used only for updating the generator during adversarial training:
@@
The full loss function, including the seed interpretation model’s loss, is described in Appendix B.
@@ 4.1 Dataset
In this experiment, we use two datasets consisting of multi-turn sentences to train the seed interpretation model and perform the text synthesis experiment.
DailyDialog [^3] http://yanran.li/dailydialog Li et al. (2017) is multi-turn conversation data used for training open-domain dialogue generation models. It consists of chit-chat-style multi-turn English conversations, and we select this data for domain-independent text synthesis. This dataset is used to evaluate the performance of TESGAN and other baselines.
IMDb [^4] https://huggingface.co/datasets/imdb Maas et al. (2011) contains highly polar movie reviews and is widely used for sentiment classification tasks. Each human-written movie review consists of several sentences, and we used this data as multi-turn data. This dataset is rougher and has a larger volume compared to DailyDialog. We evaluate the general applicability by synthesizing sentences based on IMDb-trained TESGAN. Statistics of the two datasets are shown in Appendix C.
@@ 4.2 Training Steps
First, the seed interpretation model must be pre-trained with multi-turn data to interpret the seeds. In the performance and general applicability experiments, we train the model on the 11k and 25k multi-turn sets of DailyDialog and IMDb, respectively, as shown in Table 7. Then, the model that achieves the highest BLEU-4 score in the validation set is selected in each experiment.
The second step is adversarial training. After pre-training the seed interpretation model, the generator and the discriminator learn through adversarial training. For adversarial training, real and fake seeds are created by the embedding part of the frozen seed interpretation model and the generator, respectively. Since real seeds can be generated from a significant number of sentences, all 87k and 300k sentences in each training set used in the experiment mentioned above are used to create the real seeds via Equation 3. We also generate the same number of fake seeds as real seeds for adversarial training, and the following equation represents what the discriminator and generator aim to optimize during adversarial training:
@@
where $d_{\alpha,\beta}$ means SSD, SOD respectively. $\mathcal{D}_{\mathcal{L}}$ implies updating the parameters of the discriminator to accurately predict real seeds as 1 from the perspective of real seeds. $\mathcal{G}_{\mathcal{L}}$ also means updating the generator so that the discriminator predicts the fake seeds created by the generator as 1. This approach helps partially resolve the learning imbalance problem between the generator and the discriminator Goodfellow et al. (2014). Further discussion of the above pseudocode and optimization methods is covered in Section 6.1. Lastly, hyperparameters and experiment setup are described in Appendix D.
@@ 4.3 Evaluation Metric
Target-oriented evaluation metrics, such as BLEU and ROUGE Lin (2004), are not suitable for evaluating synthetic text. This is because each synthesized sentence from random noise has no corresponding target, and the generative models aim to synthesize plausible data based on real data distribution without copying the training data. Therefore, we employ several metrics that can evaluate unconditional text generation.
@@ 4.3.1 Fréchet BERT Distance (FBD)
de Masson d’Autume et al. (2019) proposed Fréchet Embedding Distance (FED) to evaluate the quality of synthetic text, inspired by Fréchet Inception Distance (FID) Heusel et al. (2017). Alihosseini et al. (2019) proposed FBD, an improved version of FED, to measure the quality and diversity of synthesized text using a pre-trained BERT. The features of real and synthesized text obtained by the pre-trained BERT are assumed to have Gaussian distributions, and FBD is the distance between them:
@@ 4.3.2 Multi-Sets-Jaccard (MSJ)
Each synthesized sentence has no corresponding target; thus, we select MSJ Alihosseini et al. (2019), which calculates the score between real and synthesized text sets. The Jaccard Index determines the similarity of two sets, calculating the ratio of the cardinality of their intersection to that of their union. Inspired by the Jaccard Index, MSJ focuses on the similarity of the n-gram frequencies of text in the two sets, $s_{r}$ and $s_{f}$, which are the real and synthesized text sets, respectively:
$\small MSJ_{n}=\frac{\sum_{g\in G_{n}}\min\big{(}C_{n}(g,s_{r}),C_{n}(g,s_{f})\big{)}}{\sum_{g\in G_{n}}\max\big{(}C_{n}(g,s_{r}),C_{n}(g,s_{f})\big{)}}$ (13)
where $G_{n}$ and $C_{n}(g,s)$ mean the n-gram in $s_{r}\cup s_{f}$ and the normalized counts of the n-gram in set $s$. Additionally, this n-gram-based synthesized sentence evaluation method is a common approach in the field of unconditional text generation Yu et al. (2017); Press et al. (2017); Fedus et al. (2018).
@@ 4.3.3 Language Model score (LM)
de Masson d’Autume et al. (2019); Caccia et al. (2020) proposed LM, which can evaluate the quality of generated samples using a well-trained language model. LM measures the quality of generated samples, meaning that scores of the bad samples are poor under a well-trained language model. We select the pre-trained GPT-22 as a well-trained language model. LM is calculated as the cross-entropy results between the output and input of GPT-2.
@@ 4.3.4 Data Synthesis Ratio (DSR)
DSR considers not only the data memorization ratio between the training and synthesized data but also synthetic diversity itself. Short sentences identical to training data, such as ”I’m fine”, can be synthesized by coincidence. Therefore, sentences longer than two-thirds of the maximum sentence length that perfectly reproduce the training data are considered memorized data. Considering these conditions, we can calculate DSR using the following equation:
@@
where $S_{syn}$ and $S_{train}$ indicate synthesized and training text set respectively. $S_{unq}$ means the set of unique sentences of synthesized text results. If the synthesized sentences in $S_{syn}$ do not reproduce any of the sentences in $S_{train}$, $R_{syn}$ would be 1. Similarly, if the synthesized text in $S_{syn}$ is all unique, $R_{unq}$ will be 1. The final DSR is calculated as the harmonic mean of $R_{syn}$ and $R_{unq}$ ratios.
4.3.5 Self-BLEU (SBL)
Zhu et al. (2018) first proposed SBL to measure diversity of token combination. The original BLEU evaluates the degree of n-gram overlap (similarity) between one hypothesis sentence and multiple reference sentences. However, unconditionally generated text does not have specific targets, so it is not suitable for BLEU evaluation. SBL is widely used to solve this problem. SBL can evaluate n-gram-level similarity by regarding one sentence as a hypothesis and the rest as references in a synthetic text set. Since SBL evaluates based on the generated text set itself, it is not able to evaluate the quality of the synthetic text, but it is possible to evaluate the diversity of token combinations based on n-gram. Additionally, the difference between SBL and DSR lies in their evaluation criteria. DSR assesses data memorization by comparing the generated text set with the training dataset, while also considering the diversity of not n-gram-based but generated complete sentences themselves.
@@ 4.4 Baselines
In this paper, we compare our two models with the following approaches: LSTM-based Maximum Likelihood Estimation (MLE-L), PG-BLEU, SeqGAN, RankGAN, and MaliGAN. MLE-L represents the pre-training result of the generator, which all the other models undergo before adversarial training. The pre-trained generators with the lowest loss in the validation set were chosen for each method, including MLE-L. We compare these models with our original TESGAN and P-TESGAN, which is trained by adding zero-centered normal distribution noise $z$ to the generator’s output. We also evaluate the GPT-2-based pre-trained seed interpretation model (MLE-G) used in the TESGAN framework. Since MLE-based models are trained without adversarial training, they are shown as baselines in Figure 6. Finally, to demonstrate that the results of the TESGAN-based models are not solely dependent on the seed interpretation model but rather on seeds created by the generator, we present the outcomes when using Gaussian random noise as input for the seed interpretation model.
@@ 5 Results
Figure 6: Illustration showing the results of the text-GAN models. In previous research, adversarial training is conducted after the generator pre-training. MLE is represented as a baseline because it is a supervised pre-trained generator without adversarial training.
[…]
TESGAN (ours) 2.899 0.042 0.021 0.967 (1, 0.936) 4.236 0.743 0.623
P-TESGAN (ours) 2.274 0.032 0.014 0.841 (0.997, 0.727) 3.642 0.790 0.702
SeqGAN Yu et al. (2017) 6.153 0.040 0.015 0.880 (0.883, 0.877) 5.094 0.420 0.266
RankGAN Lin et al. (2017) 6.409 0.048 0.023 0.890 (0.895, 0.886) 5.123 0.446 0.290
MaliGAN Che et al. (2017) 21.436 0.003 0 0.030 (1, 0.015) - - -
PG-BLEU Yu et al. (2017) 9.002 0.015 0.006 0.569 (0.555, 0.584) 4.584 0.628 0.484
MLE-L Yu et al. (2017) 6.284 0.045 0.021 0.955 (0.925, 0.987) 5.168 0.403 0.242
MLE-G 9.592 0.026 0.014 0.078 (1, 0.040) 3.543 0.948 0.944
Random Noise $\dagger$ 14.142 0.016 0.006 0.930 (1, 0.869) 4.562 0.516 0.404
Table 1: Performance of the models. P-TESGAN denotes the perturbed TESGAN. $\dagger$ is the result of directly entering Gaussian random noise as an input to the seed interpretation model. The second group of models consists of autoregressive models. $\ast$ denotes a metric not considered when selecting the best model. - denotes that the confidence of the result is low because the quality of the synthesized sentence is poor. The superscript of each metric represents what each metric can measure (Q: quality, D: diversity, M: data memorization).
TESGAN (17-epoch, DailyDialog) P-TESGAN (10-epoch, DailyDialog) Random Noise
@@
TESGAN (18-epoch, IMDb)
This is probably one of the best of the best of the series.
I was bored to think about how stupid this movie was.
”The Deadly Loved One” is the story of a rebellious college basketball
I have to say, this is the worst film I have ever seen.
I was very excited to see it, anticipating Christmas eve.
This movie was one of the best of the year for me.
Table 2: Example of unconditionally synthesized sentences. P-TESGAN denotes the perturbed TESGAN.
5.1 Metric-based Evaluation
In this section, we compare the results of the synthesized text at every epoch of adversarial training using the metrics mentioned in Section 4.3. This experiment was performed with models trained on the DailyDialog dataset. Since Fréchet BERT Distance (FBD) and Multi-Sets-Jaccard (MSJ) require a real text corpus, the test set is used as the real text corpus. Data Synthesis Ratio (DSR) is calculated with the training set as the data memorization ratio needs to be computed.